Circular No. 5681

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA CYGNI 1992 T. Hayward, Cornell University, writes: “Recent observations have

revealed new features in the 7.5- to 13-μm spectrum of N Cyg 1992. Spectrograms obtained with the Mt. Lemmon 1.5-m telescope (+ HIFOGS) by D. H. Wooden and F. C. Witteborn (NASA Ames Research Center), and A. L. Sprague (Lunar and Planetary Laboratory, University of Arizona) on Oct. 17 and Nov. 8 (days 243 and 263 after the eruption) show [Ne VI] 7.65-μm and [Ar III] 8.99-μm lines, as well as the [Ne II] 12.81-μm line reported on IAUC 5497 and 5537. Low-resolution spectrograms obtained by T. Hayward, J. Houck, and J. Miles (Cornell University) using the 5-m Hale telescope (+ SpectroCam-10 = SC-10) on Nov. 14-15 (days 269-270) confirm all three lines. In addition, SC-10 spectra at resolution 150 km/s resolve each line at FWHM ∼ 2200 km/s, although about one-third of the [Ne VI] line is totally obscured by a deep telluric absorption feature at 7.66 μm. The following integrated intensities are given in units of 10^{−12} erg s^{−1} cm^{−2}: [Ne VI], 218 ± 9 on Oct. 17 and Nov. 8, 266 ± 13 on Nov. 14-15; [Ar III], 8 ± 3, 10.3 ± 0.6; [Ne II], 5 ± 1, 7.0 ± 0.3. The [Ar III] line is not present in similar SC-10 data taken on Aug. 9, but we cannot rule out [Ne VI] emission on that date because of poor atmospheric transmission that would have obscured the line. The free-free continuum in the November SC-10 spectrum was 0.18 ± 0.04 Jy at 10.5 μm, but the Humphreys α line at 12.37 μm reported by the Cornell group on IAUC 5497 and reobserved in August was not detected. The appearance of the [Ne VI] and [Ar III] lines confirm that the nova has entered a coronal line phase, and these are the first observations of the [Ar III] line in a nova.”

Visual magnitude estimates: Nov. 14.72 UT, 9.5 (H. Dahle, Oslo, Norway); 18.79, 9.8 (M. Tombelli, Montelupo, Italy); 30.29, 10.6 (A. Mizser, Budapest, Hungary); Dec. 12.06, 9.8 (B. H. Granslo, Fjellhamar, Norway); 15.83, 9.9 (P. Schmeer, Bischmisheim, Germany); 17.70, 9.9 (Granslo).

PERIODIC COMET SWIFT-TUTTLE (1992t) Total visual magnitude estimates: Dec. 11.04 UT, 5.0 (R. Keen, Mt.

Thorodin, CO, 8×40 binoculars); 13.71, 5.4 (R. Haver, Tolfa, Italy, 10×50 binoculars; 2^{o} tail in p.a. 55^{o}); 15.72, 5.2 (P. Schmeer, Bischmisheim, Germany, 20×80 binoculars); 16.70, 5.4 (M. Tombelli, Montelupo, Italy, 11×80 binoculars); 18.95, 5.2 (D. W. E. Green, Cambridge, MA, 20×80 binoculars); 20.95, 5.2 (J. E. Bortle, Stormville, NY, 10×50 binoculars).

1992 December 21 Daniel W. E. Green